OpenStudy (anonymous):

Are the complex fractions the quantity (x^2 - x - 20 / 4) / ( x - 5 / 10) and the quantity (x^2 - x - 20 / x - 5) / (4/10) equivalent? Simplify each, and then explain why or why not.

OpenStudy (anonymous):

yes both are equivalent \[(a/b)\div(c/d)=(a/b)\times(d/c)\]

OpenStudy (anonymous):

exactly as @syam_krisna , says...

OpenStudy (anonymous):

why r dey equivalent?

OpenStudy (anonymous):

try simplify both fractions without multiplying... just leave it in factored form..

OpenStudy (anonymous):

ok.. let a = x^2 - x -20 b = 4 c = x-5 d = 10

OpenStudy (anonymous):

that first quantity is (a/b)/(c/d) correct?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

let's simplify this.. \[\large \frac{a}{b}\div \frac{c}{d}=\frac{a}{b}\times \frac{d}{c}=\frac {ad}{bc}\]

OpenStudy (anonymous):

kkkkkk i got itttt hahahaha

OpenStudy (anonymous):

now do that with theother...

OpenStudy (anonymous):

at least i got practice my LaTex ing...

OpenStudy (anonymous):

the first one gives me x+4/40 right

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

u sure???????? wat u got?

OpenStudy (anonymous):

5(x+4)/2

OpenStudy (anonymous):

want me to show?

OpenStudy (anonymous):

u sure? the equation is (x^2 - x - 20) / (4) / ( x - 5) / (10) try it again the problem is wrote the equation wrong

OpenStudy (anonymous):